OpenStudy (justcatx3):

Sketch the graph of h(x)=x^2-3x-4/x-4. Label any important features. State the domain and range.

OpenStudy (justcatx3):

Ok, so I already figured out the domain. Set x - 4 to equal zero. x - 4 = 0 x - 4 + 4 = 0 + 4 x = 4 {x|x =/= 4} However, I'm unsure how to find the range. I also put the equation in a graphing calculator and found the line to be straight. So I'm not sure what label any important features might mean. I know it doesn't have any extremem values since the line is straight.

jhonyy9 (jhonyy9):

x^2 -3x-4 = (x-4)(x+1)

OpenStudy (justcatx3):

Slightly confused xD Are you saying I should factor it?

OpenStudy (justcatx3):

@thesmartone

OpenStudy (mathmale):

Yes, jhonny is suggesting you factor the numerator. If you do that, you see that x-4 in numerator and denominator can be cancelled out. Even so, this does NOT mean that x=4 is an acceptable input; it is not.

OpenStudy (justcatx3):

Thank you for explaining~ So that would leave me with x+1. What is this helping me find though? The range?